OpenStudy (anonymous):

Medal! Bonnie has a container in the shape of a rectangular pyramid. The formula for the surface area of the enclosed space is S = lw + 0.5Ph. Solve for P. P = S – lw – 0.5h P = S + lw + 0.5h P = The quantity S minus l times w all divided by 0.5 times h P = S divided by the quantity l times w plus 0.5 times h

KjSaif (kjsaif):

C

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

??? where your text went

OpenStudy (anonymous):

Firstly, Let's subtract lw from both sides as follows: S=lw + 0.5 Ph

OpenStudy (anonymous):

s (-lw) = lw(-lw) + 0.5 Ph this gives: s-lw = 0.5 Ph

OpenStudy (anonymous):

Next, Ph really means P x h , and all we want is P, so we will need to divide both sides by h, as follows: s-lw/h = (0.5 Ph)/h

OpenStudy (anonymous):

now the h's on the right-hand side cancel, and you are left with: s-lw/h = 0.5 P to get P that means we need to multiply both sides by 2 as follows: 2 x(s-lw/h) = 2 x 0.5 P

OpenStudy (anonymous):

so our answer is 2(s-lw/h) = P or (2s -2lw)/h = P